OpenStudy (ilovemath14):

HELP ASAP!! DUE TOMORROW!! FAN AND MEDAL!! 6. 3,700 deposit earning 3.3% compounded monthly, after 5 years. The balance after 5 years will be $-------- 8. Suppose the atmospheric air pressure at sea level is about 105 kilopascals. For every, 1000-m increase in altitude, the pressure decreases about 11.5%. What's the approximate pressure at an altitude of 3000 m?

OpenStudy (jdoe0001):

hmmm anyhow, montly, means, 12 times in a year, thus, the cycle is 12 = n thus \(\qquad \textit{Compound Interest Earned Amount}\\ A=P\left(1+\frac{r}{n}\right)^{nt} \qquad \begin{cases} A=\textit{current amount}\\ P=\textit{original amount deposited}\to &\$3,700\\ r=rate\to 3.3\%\to \frac{3.3}{100}\to &0.033\\ n=\textit{times it compounds per year}\to &12\\ t=years\to &5 \end{cases}\)

OpenStudy (ilovemath14):

@TheSmartOne @confluxepic @qhhelper @KendrickLamar2014 @jabez177 @Preetha

OpenStudy (ilovemath14):

3700 (1+0.033/12)^5 @jdoe0001